OpenStudy (anonymous):

Will fan you!! Write each expression using the base only A. 9^-8 x 9^-2 x 9^10 B. 5^4 x 5^7 C. 5^-5 x 5^7

OpenStudy (anonymous):

when adding exponents i get 9^0 ....

OpenStudy (anonymous):

really it is 9^10 / (9^8 * 9^2).... i took the negative exponents and put them in the denominator ...

OpenStudy (anonymous):

i dont get it :/

OpenStudy (anonymous):

any negtive exponent say 9^-8 is the same as writing 1/9^8

OpenStudy (anonymous):

Alright i get that how do you get am answer though

OpenStudy (anonymous):

simplifying it should be 9^10/ 9^10 which cancels 1/1 = 1 but can be written as 9^0 = 1

OpenStudy (anonymous):

alright so B. would be 5^11?

OpenStudy (anonymous):

yeah that's what i got

OpenStudy (anonymous):

if u had (5^4)^3 u would mulitply them so it would be 5^12

OpenStudy (anonymous):

sweet! and C im haveing a hard time with that one

OpenStudy (anonymous):

if ur asked that (5^4)^3 another way is u can think of it as 5^4 *5^4 *5^4 then u rewrote it like the others so u can just add them now and u get the same answer 5^12

OpenStudy (anonymous):

yeah i get that, but what about 5^-5 x 5^7

OpenStudy (anonymous):

the expression can be rewritten as 5^7/5^5

OpenStudy (anonymous):

so it's like 5 * 5 * 5 * 5 * 5 *5 *5 / 5 * 5 * 5 * 5 *5 so u cancel5 of them and u get 5^2 /1 so 5^ 2 if simplified would be 25

OpenStudy (anonymous):

SO pretty much leave 5 alone then add^-5 and ^7 youd get ^2 then you have 5^2 then you add that?

OpenStudy (anonymous):

yes, but the 5^2 is 5 * 5 = 25

OpenStudy (anonymous):

when the problems deal with more variables then it's a lot easier to make a mistake if not rewritten and worked out in serveral steps

OpenStudy (anonymous):

Yeah thats what i meant ha thanks!

OpenStudy (anonymous):

it's cool glad i could help.

OpenStudy (anonymous):

Wanna help a little more?

OpenStudy (anonymous):

well i can try but i don't know it all

OpenStudy (anonymous):

Its ok its pretty basic

OpenStudy (anonymous):

A. (y^-3)^6 B. (-h^4)^5 C. (t^-4)^-9t^2 D. (5t^3)^-4

OpenStudy (anonymous):

so( y^3)^6 is same thing as y ^3*6 = y^18

OpenStudy (anonymous):

ok i see how you did that so B would be -h^20?

OpenStudy (anonymous):

yes but if it were (-h^4)^4 it would be positive h^16

OpenStudy (anonymous):

-1^5 remains negative lot of mistakes are here

OpenStudy (anonymous):

so -1^4 would be positive

OpenStudy (anonymous):

Alright

OpenStudy (anonymous):

is question c written like that? (t^-4)^-9t^2?

OpenStudy (anonymous):

yeah it is

OpenStudy (anonymous):

so on part d can be rewrttten as 5^-4 * t^3*-4 then rewrtten as 1/ 5^4 t^12

OpenStudy (anonymous):

can be simplified 1 / 625t^12

OpenStudy (anonymous):

i actually get that haha

OpenStudy (anonymous):

we didn't do C

OpenStudy (anonymous):

so t^-4*-9t^2 = t^36t^2 i think that is the answer

OpenStudy (anonymous):

abc

OpenStudy (anonymous):

sorry